A 12.0kg box is being pulled alone level ground at constant veloncity by a horizontal force of 38.0 newton .What is the cofficient of kinetic friction between the box and the floor?


1
Expert's answer
2021-05-17T17:35:38-0400

Let's apply Newton's Second Law of Motion:


"F_{pull}-F_{fr}=0,""F_{pull}=F_{fr}=\mu_k mg,""\mu_k=\dfrac{F_{pull}}{mg}=\dfrac{38\ N}{12\ kg\cdot9.8\ \dfrac{m}{s^2}}=0.32"